2. You own a small farm near a large city, and you are about to decide whether to work on that small farm or take a job in the city. Your utility depends on your income per day (Y) and on hours of leisure (L). Your daily income from farm work is:


Yf = 20hj - 0.2h2f


where hr is hours of work on the farm, and your daily income from the city job is:


Ye = 1 4hc


where he is hours of work in the city.


a. If you can work either on the farm or in the city, but not both, which sector would you choose?

b. If you can work both on the farm and in the city, how would you allocate your time?




1
Expert's answer
2022-01-25T08:44:42-0500

a) I would choose to work in the city, considering the income being higher than working in the farm

"Yc=14hc"

"hc=8hours"

"Yc=14(8)=112"

b)"Yf=20hf-hf^{2}"

"Yf=20(8)-8^{2}=96"

If the income is 96 for "Yf" in 8hours

1hour=12

For "Yc" 1 hour= 14

I would allocate more time to work in the city than in the farm

Since I have 8hours each day to dedicate to work, I would allocate 5hours to work in the city and the remaining 3hours to work in the farm.